### v1

# Adaptive Interactive Navigation of Quadruped Robots using Large Language Models

Abstract
Robotic navigation in complex environments remains a critical research challenge. Traditional navigation methods focus on optimal trajectory generation within free space, struggling in environments lacking viable paths to the goal, such as disaster zones or cluttered warehouses. To address this gap, we propose an adaptive interactive navigation approach that proactively interacts with environments to create feasible paths to reach originally unavailable goals. Specifically, we present a primitive tree for task planning with large language models (LLMs), facilitating effective reasoning to determine interaction objects and sequences. To ensure robust subtask execution, we adopt reinforcement learning to pre-train a comprehensive skill library containing versatile locomotion and interaction behaviors for motion planning. Furthermore, we introduce an adaptive replanning method featuring two LLM-based modules: an advisor serving as a flexible replanning trigger and an arborist for autonomous plan adjustment. Integrated with the tree structure, the replanning mechanism allows for convenient node addition and pruning, enabling rapid plan modification in unknown environments. Comprehensive simulations and experiments have demonstrated our method’s effectiveness and adaptivity in diverse scenarios. The supplementary video is available at page: https://youtu.be/W5ttPnSap2g.
Index Terms:
Interactive navigation, large language model, task and motion planning, quadruped robot.
I Introduction
Navigation in diverse and complex environments is a pivotal challenge in robotics, necessitating innovative approaches to ensure effective and adaptable decision-making and planning. Recent research has demonstrated the potential of robotic platforms to traverse cluttered environments and navigate to the desired position, such as utilizing quadrupedal robots [1, 2] for maneuvering across various terrains. Despite the progresses, conventional navigation strategies primarily focus on collision-free trajectory planning within a given free workspace, which significantly limits their applicability in real-world scenarios such as disaster rescue or warehouse logistics, where collision-free paths may not exist, and the robot must interact with the environment to create viable routes.
To address these limitations, previous works have explored navigation among movable obstacles (NAMO), where the robot can actively interact with and reposition movable obstacles to create viable paths. This NAMO research expands the scope of robotic navigation, enabling operation in highly cluttered and confined environments. However, some prior research [3] focused on global planning within movable obstacles but idealized interactive behavior in simplified simulations, which limited their applicability in real-world situations. To address this limitation, Simon et al. [4] have explored integrating mobile platforms with robotic arms, employing whole-body control to relocate obstacles and create pathways, and modeling more realistic interaction behaviors to enhance applicability in real-world environments. However, NAMO approaches only focus on removing obstacles without utilization of movable obstacles as tools, such as leveraging shorter obstacles as stepping stones to help the robot traverse higher hurdles impassable originally, thereby limiting their abilities to solve long-horizon interactive navigation tasks that require multi-step strategic utilization of environmental objects.
Figure 1: Interactive navigation in challenging environments. In contrast to conventional obstacle-avoidance navigation systems that passively adapt to predefined free space and fail to reach goals in environments without feasible paths, our method (AINav) can push aside obstacles to create pathways within congested obstacles or utilize tools to traverse excessively high hurdles, tackling interactive navigation tasks in challenging environments.
Leveraging the advanced understanding and reasoning abilities of large language models (LLMs), recent works have utilized LLMs combined with pre-trained atomic skill libraries to interact with environments to address complex long-horizon tasks [5, 6]. They formulate the complex interactive tasks as task and motion planning (TAMP) problems, where task planning involves task decomposition to assist in completing final tasks, such as selecting which objects to interact with and determining the sequence of interactions, while motion planning considers how to execute the high-level task by a controller to accomplish the task. However, these methods require global scene description from an omniscient perspective, which is often unavailable in real-world scenarios where robots operate with limited, egocentric observations. This limitation restricts their ability to adapt to dynamic and partially observable environments. Closed-loop task planning can respond to new observations through iterative replanning [7, 8]. Nevertheless, previous works commonly focus on passive failure recovery, lacking the ability to proactively utilize new information and adjust plans, which represents a critical aspect for effective navigation in partially observable environments. The primary challenge we aim to address is how to create effective plans for navigation tasks in unknown interactive environments while adaptively responding to newly acquired information.
To overcome these limitations and challenges, we propose the Adaptive Interactive Navigation (AINav) approach to tackle interactive navigation tasks (Fig. 1). The proposed system demonstrates dual capabilities: it actively interacts with the environment to change the workspace for achieving navigation objectives, and adaptively reacts to new observations in unknown environments. This dual functionality enables effective navigation in unfamiliar and interactive environments using egocentric sensing information, without relying on global environmental knowledge. While recent advancements in robot parkour have demonstrated the capability to traverse across varying heights relative to the size of the quadruped robot, our research emphasizes enabling the robot to actively manipulate the environment, tackling challenges that extend beyond the scope of conventional passive obstacle avoidance and terrain adaptation navigation systems. The main contributions can be summarized as follows:
  • We propose a primitive skill tree for task planning using LLM, which represents task decomposition as a tree structure with atomic skills. This design enables robust reasoning by exploration of multiple potential solutions and rapid adaptation to new information by tree structure modification.
  • We develop a comprehensive skill library with reinforcement learning (RL) to equip the robot with advanced locomotion and interaction capabilities. Integrated with customized reward design and task curriculum, robots can learn to navigate and manipulate robustly in complex environments.
  • We introduce an adaptive replanning mechanism including two LLM-based modules, an advisor and an arborist, where the advisor adaptively triggers replanning based on new observations, while the arborist automatically reconstructs the decision tree for optimal task execution.
We validate our method across comprehensive simulations and real-world experiments, revealing its efficacy in tackling challenging navigation tasks and its capacity for fast adaptation to new observations in unknown environments. Through this novel framework, we aim to extend the boundaries of robotic navigation, offering an effective and computationally efficient solution for interactive navigation in complex environments. The remainder of this paper is organized as follows: Section II reviews related work in robotic navigation and interactive planning. Section III presents the proposed AINav framework with specific methods. Section IV describes the experimental setup and results. Finally, Section V concludes the paper and discusses future directions.
II Related Work
II-A Quadruped Robot Navigation
Quadruped robots have demonstrated agile capabilities and terrain traversability in diverse complex environments in recent researches [9, 10]. Based on the superior agility and adaptivity, quadruped robot navigation has made significant progress through various kinds of approaches, including sampling-based method [11], learning-based method [12, 13]. Recently, benefiting from the powerful understanding and reasoning ability of pre-trained large models, abundant works utilized the LLM and vision-language models (VLM) to understand the semantic environment information for more flexible terrain traverse, so as to tackle navigation tasks [14, 15]. However, current approaches often assume the existence of navigation paths, neglecting the possibility that such paths may not exist in real cluttered scenarios. Zhang et al. [16] presented an interactive navigation framework that instructs robots to navigate in an environment with traversable obstacles with LLMs. However, they aim to adapt to the environment passively rather than actively utilizing interactive objects for task completion. Wu et al. [5] and Xu et al. [6] proposed using LLMs to interact with environments and employ tools to address complex long-horizon tasks with quadruped robots more effectively. However, they assume complete scene descriptions with omniscient information and perform open-loop planning, which is impractical for navigation tasks where the robot obtains real-time egocentric observation incrementally.
II-B Navigation Among Movable Obstacles
Navigation among movable obstacles refers to the problem of robot navigation that requires manipulation of the environment. Early research [3] mainly concentrated on modeling abstract and idealized interactive behaviors, overlooking the complexities inherent in real-world physical interactions. While these approaches are effective in simplified simulation settings, their failure to account for real-world physics significantly restricts their practical applicability. Recent work by Simon et al. [4] has advanced the field by integrating mobile platforms with robotic arms and utilizing whole-body control to relocate obstacles and clear pathways. This approach incorporates more realistic interaction models, improving applicability in real-world environments. Nevertheless, current NAMO approaches predominantly focus on obstacle removal and do not fully capitalize on the potential of movable obstacles, such as utilizing shorter obstacles as steps to traverse higher barriers. This limitation restricts their ability to address more complex long-horizon tasks in more intricate scenarios.
II-C Closed-Loop Task Planning with LLMs
Due to superior performance in language understanding and commonsense reasoning, LLMs are widely adopted to efficiently solve closed-loop task planning, which refers to generating a series of intermediate steps to achieve the specific goal while modifying the plan based on new observations. Many works incorporate an iterative manner for closed-loop task planning, which replans based on the description or analysis of real-time observations [17, 18]. However, these methods typically generate a single plan for execution, which often contains errors or infeasible decomposition due to model uncertainties. This increases the need for replanning and hinders computationally efficient planning. Tree-based structure for LLM reasoning and planning enhances the solution quality by generating multiple plans and select the optimal tree path with state evaluation [19, 7, 20]. However, these methods either lack a replanning mechanism or trigger replanning solely upon encountering failures and concentrate primarily on error correction, overlooking a comprehensive understanding and effective utilization of interactive environments.
Figure 2: An overview of AINav approach, a hierarchical system for interactive navigation. The task planning module processes visual input to generate a subtask skeleton for execution by the motion planning module, integrated with an adaptive replanning mechanism that enables flexible replan triggering and rapid adjustments to the task plan.
III Adaptive Interactive Navigation
Our framework consists of three primary components: LLM-based task planning, motion planning with pre-trained skills, and adaptive replanning, as shown in Fig. 2. The robot first captures the first-person perspective images and extracts scene understanding and object descriptions using a VLM and object detector, respectively. By combining user instructions with environmental information as inputs, the task planning module uses the LLM to construct a primitive tree for reasonable task decomposition and convenient adjustments based on new information, enabling the robot to obtain the optimal high-level primitive skeleton. Leveraging a pre-trained skill library consisting of robust locomotion and interaction skills with RL, the motion planning module executes specific low-level actions following the skeleton to complete the task. Finally, a novel adaptive replanning mechanism interprets new observations from an egocentric perspective to determine whether replanning is necessary, allowing for swift reaction to incremental environmental observations.
III-A Task Planning with Primitive Skill Tree
We propose the primitive skill tree for task planning. The primitive skill tree encompasses two LLM-based roles: a proposer for primitive tree construction, and an evaluator for node evaluation and skeleton selection.
Proposer: object-level to skill-level. Given the user instructions $I$ and environmental information $O$, we first utilize the LLM to propose multiple potential object-level planning strategies $P_{obj}$, where each plan specifies which objects to use and how to leverage them to achieve the task
$\begin{split}P_{obj}=\text{LLM}(I,O).\end{split}$ (1)
Building upon object-level plans, we subsequently generate standardized skill-level plans $P_{skill}$, where each step corresponds to a primitive skill from the robot’s skill library $S$.
$\begin{split}P_{skill}=\text{LLM}(P_{obj},S).\end{split}$ (2)
After skill-level plans are generated, we construct the primitive skill tree by merging steps of plans in $P_{skill}$ with identical historical traces. This approach offers two key advantages: First, unlike other approaches that use the LLM to directly generate a single task plan, multiple alternative plans proposed by the LLM are more robust to unexpected errors, such as misunderstandings by the LLM, limited scene observation, and environmental uncertainties. Second, compared to directly using the LLM to generate detailed plans that may lead to illogical or inexecutable plans, transitioning from holistic object utilization to specific skill planning tends to propose more reasonable and feasible object interaction strategies.
Evaluator: short-term to long-term. To evaluate the proposed plans, we utilize another LLM as an evaluator to assess the immediate reward $r(n)$ of each node $n$ in the primitive tree, which considers the difficulty of primitive skills and the contribution to the specific task. For example, if the robot interacts with objects like pushing movable obstacles to create a free path for navigation, this skill is essential to completing the task, but the interaction is also difficult to complete. If one skill is non-executable, the subsequent nodes including this node will be deleted to ensure the feasibility of the plan. If a trajectory in the primitive tree reaches the target point at the end, an additional bonus is given to the leaf node as the terminal reward. Finally, we apply a backup method to update the long-term cumulative reward $Q(n)$ of nodes with
$\begin{split}Q(n)=r(n)+\gamma\cdot\frac{1}{|\text{children}(n)|}\sum_{n_{c}\in\text{children}(n)}Q(n_{c}),\end{split}$ (3)
where $\gamma$ is the discounted factor. By iteratively selecting the child nodes with maximal cumulative reward, we derive the optimal high-level primitive skeleton. This approach also allows for the existence of alternative strategies, enabling rapid replanning upon receiving new environmental information, which will be specifically addressed in Section III-C.
III-B Skill Pre-training with Reinforcement Learning
To robustly execute the decomposed tasks, we train a primitive skill library using RL for motion planning. The skill library includes both low-level locomotion skills, such as walking and climbing, and high-level skills like navigation and object-pushing strategies. We first train locomotion policies to develop fundamental movement capabilities on diverse terrains. Building upon the locomotion policies, we employ a hierarchical RL framework to train more advanced strategies, such as navigation policy for collision-free movement in cluttered environments and pushing policy for moving objects to designated positions.
In low-level locomotion skill training, the walking policy training is formulated as a velocity-tracking task, which encourages robots to follow the linear and angular velocity command while adapting to arbitrary terrain. As for the climbing skill, to ensure the robot learns an effective policy, we refine the reward design to encourage the robot to reach the goal while aligning its movement direction toward the goal. Specifically, we add a customized head collision penalty to prevent reliance on the head when climbing up high hurdles, encouraging the robot to use its legs as support instead of the head to learn a more natural and safe climbing action. Additionally, curriculum learning is employed to gradually increase terrain difficulty or hurdle height, enabling the development of robust locomotion policies.
With low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow. Specifically, the pushing skill aims to push boxes to the target pose, and the navigation skill navigates the robot to the target pose while avoiding obstacles. We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill learning, facilitating robust high-level skill learning. All these skills are trained using PPO [21] in the IsaacLab simulation environment. After training these primitive skills, we encapsulate them into APIs for task planning, enabling sequential skill execution to accomplish complex tasks.
III-C Adaptive Replanning with Observation Interpretation
Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot continuously acquires new observations with an egocentric perspective. For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan based on new environmental comprehension. This mechanism consists of two LLM-based roles: an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
The advisor analyzes new environmental information from visual perception and the current plan to determine whether replanning is necessary based on these comprehension updates. We let the advisor conduct an environmental interpretation in three aspects, including failure, new objects, and revaluation. Failure indicates that the current plan has encountered some errors, new objects represent the discovery of a new interactive object that may be useful, and revaluation means reassessing the plan based on updated information about the environment. Unlike previous works that only replan based on failure, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment. For example, when you discover a new object, you need to analyze its geometric and semantic information and decide to utilize it if it is more helpful to your task. As for revaluation, the robot’s understanding of the object information under observation uncertainties will be updated, which will affect the evaluation of the current plan. Based on these three types of comprehension, the advisor will decide whether to replan and replan suggestions.
When the advisor determines that replanning is necessary, it generates specific suggestions for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations. Node addition incorporates new object information, while node pruning removes failed paths. This tree-based architecture enables computationally efficient replanning through its inherent structural flexibility, facilitating real-time task execution. Following structural modifications, the system performs backup operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills. Through the adaptive replanning approach, the robot can respond to new environmental information flexibly and adjust the planning strategies accordingly. By analyzing new observations, the robot can assess their impact on task completion and decide whether adjustments to the current plan are required. This process helps avoid the computational cost of unnecessary replanning.
IV Simulation
In this section, we present the simulation experiments conducted to validate the effectiveness of the proposed method. The experiments are structured into four main parts: (1) the experimental setup, which outlines the simulation environment and interactive navigation tasks; (2) the baseline methods compared to AINav to establish a performance benchmark; (3) the simulation results and analysis; (4) ablation studies to investigate the contribution of individual components of the proposed method.
IV-A Experimental Setup
Simulation Settings. To validate the effectiveness of our proposed method, we conduct extensive simulations in IsaacLab. The simulations are conducted in a $8$m $\times$ $8$m space with fixed obstacles and movable boxes. We design diverse simulation environments corresponding to four interactive navigation tasks, which will be detailed in the following section. We use the Unitree Go2 as the robotic platform to perform the simulations. All experiments are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
Task Description. We design four interactive navigation tasks: box obstruction, box usage, stair building, and an integrated task, as illustrated in Fig. 3. The box obstruction task involves environments where boxes block the path. Box usage and stair building require the robot to interact with boxes to complete navigation tasks. The integrated task combines challenges from the previous three tasks. Each scenario is divided into three levels of difficulty: low, medium, and high. The low-difficulty scenario can be regarded as the traditional navigation task where the goal point can be reached without interaction. Medium difficulty requires an interactive environment, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with immovable boxes, designed to evaluate the replanning performance and efficiency of various methods. The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters. We set the maximum simulation time to $120$s and conducted $10$ repeated experiments for each scenario.
Evaluation Metrics. The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), and execution time (ET) to comprehensively assess the performance of different methods. SR refers to the probability of the robot successfully reaching the target point across 10 trials. OT denotes the total simulation time across all scenarios. If a trial fails, OT is recorded as the maximum simulation time, which is $120$s. OTS represents the overall completion time for successful task completion. PT and ET indicate the time spent on task planning and skill execution under successful completion, respectively. SR, OT, and OTS indicate the comprehensive effectiveness and efficiency of methods, while planning time and execution time reflect the computational efficiency and quality of task planning.
Figure 3: Interactive tasks in simulations.
IV-B Baseline Methods
We compare the performance of the proposed method to the following baselines. Art-planner [11] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks. Hierarchical RL trains a high-level policy to function as the task planning module, which decides which low-level skill to use and the corresponding parameters. The low-level skills utilized in hierarchical RL are identical to those in our method. The reward is inversely related to the distance between the robot and the goal, encouraging the robot to move toward the target. Furthermore, we also select two LLM-based method for comparison. The first is RoboTool [6], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges. Since this method uses open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output. The second method is Tree-planner [7], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems. Because its plan sampling module requires global information, we include the initial object information as input. We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks. For a fair comparison, all methods use the same set of low-level skills as our method.
IV-C Results and Analysis
Simulation results are displayed in Table I. We can observe that our method exhibits higher success rates and less overall time compared to other baselines. The Art-planner achieves high task success rates in low-difficulty scenarios with low execution time, showcasing its effective planning capabilities in traditional navigation tasks. However, in medium-difficulty and high-difficulty scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks. Hierarchical RL can achieve limited success in low-difficulty scenarios of box obstruction, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
TABLE I: Performance Metrics Across Different Scenarios and Methods
Method Metric Box obstruction Box usage Stair building Integrated task
L M H L M H L M H L M H
Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 80% 0% 0%
OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 57.2 120.0 120.0
OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - 41.5 - -
PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - 6.7 - -
ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - 34.8 - -
Hierarchical RL SR 40% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
OT 87.4 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0 120.0
OTS 38.5 - - - - - - - - - - -
PT 0.1 - - - - - - - - - - -
ET 38.4 - - - - - - - - - - -
Robotool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 60% 30% 0%
OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 88.0 108.7 120.0
OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 66.6 82.2 -
PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 25.7 -
ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 43.1 56.5 -
Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 70% 30% 0%
OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 90.3 116.6 120.0
OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 77.6 108.5 -
PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 29.3 36.2 -
ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 48.3 72.3 -
Ours SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 80% 60% 50%
OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 69.7 98.9 113.8
OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 57.1 84.9 107.5
PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 20.4 34.1 39.2
ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 36.7 50.8 68.3
  • L = Low difficulty, M = Medium difficulty, H = High difficulty
  • SR = Success Rate, OT = Overall time, PT = Planning Time, ET = Execution Time
Figure 4: Average success rate and overall time across all scenarios of different methods.
Figure 5: Success cases of the proposed method in four interactive navigation tasks. The red phases indicate the robot interacts with objects, while the green phases represent the robot performing replanning based on new environmental observations.
Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, Robotool and tree-planner achieve a higher success rate in interactive navigation tasks compared to traditional navigation methods. However, Robotool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks. While tree-planner leverages a closed-loop design to adjust its decisions based on new observations, its performance is highly dependent on the quality of the sampled plans, resulting in instability in planning results. Specifically, once the tree structure is generated, its inflexibility can hinder the planner from obtaining the optimal solution that does not exist in the original tree or changing to an updated better strategy in response to new observations. Additionally, both methods require omniscient environmental information to generate task decomposition plans, rendering them ineffective in unknown environments. In contrast, our method not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information. This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios, especially in challenging situations that require replanning. Although node evaluation incurs additional planning time, the improved task planning quality and flexible replanning mechanism enable our method to achieve higher success rates and shorter overall completion time compared to other LLM-based approaches. As shown in Fig. 4, AINav demonstrates the highest success rate and the lowest overall completion time across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
We also present successful cases of our method across four tasks in Fig. 5. In the box obstruction scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position. In the box usage and stair building scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which is unreachable for a quadruped robot under normal circumstances. In integrated task scenario, when confronted with various potential interaction tasks, the robot understands the current issue by VLM and intelligently interacts with the appropriate objects to facilitate task completion. Furthermore, we can observe that the robot is capable of responding to updated information, such as when the box fails to move or the robot identifies new objects and tasks. By leveraging an efficient closed-loop system, the proposed approach enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
IV-D Ablation Studies
Figure 6: Comparison results in ablation study.
To further investigate the effectiveness of the proposed method, we conduct ablation study to compare AINav with the following four variants. 1) AINav-Single, which only generates a single plan with LLM in the task planning module; 2) AINav-Skill, which directly generates skills without incorporating object-centric reasoning; 3) AINav-No, which removes the replanning module; 4) AINav-F, which only replans when a failure is encountered. We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building, which need interaction for task completion, to compare the success rate and overall time. The results of ablation studies are presented in Fig. 6. AINav-Single exhibits low success rates, as the single plan generated by the LLM is often flawed or inefficient due to inherent uncertainties of LLM, leading to task failures. Benefiting from the tree structure that enhances the robustness and planning quality, AINav-Skill achieves a reasonable success rate. However, it lacks a comprehensive understanding of object utilization in complex scenarios, leading to task decomposition plans inexecutable. AINav-No performs poorly because it lacks a replanning phase, making it unable to adapt to new observations. Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, often resulting in excessive time spent and task timeouts. By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving higher success rates and improved time efficiency.
IV-E Performance of Primitive Skills
(a) Climbing up/down skill
(b) Navigation skill
(c) Pushing skill: small rectangle box
(d) Pushing skill: big square box
Figure 7: Display of diverse RL-driven skills. (a) and (b) show the climbing and navigation skills adapted to various terrains, while (c) and (d) showcase robust interaction capabilities with boxes of different shapes and sizes.
We visualize the performance of various skills in Fig. 7. In Fig. 7(a), the robot is depicted climbing up and down a 0.3-meter height, effectively adapting to different terrains to accomplish navigation tasks. Fig. 7(b) illustrates the robot’s capability to navigate to a target point while avoiding obstacles in a cluttered environment. In Fig. 7(c) and Fig. 7(d), we showcase the robot pushes boxes of varying sizes and shapes to the designated positions, demonstrating the generalization capability of the pushing strategy. The box-pushing process typically consists of three steps: first, the robot approaches the box and positions itself so that the box is between the robot and the target position, preparing for the subsequent pushing action; second, the robot pushes the box toward the target point; and finally, the robot fine-tunes the box’s position to complete the task. These versatile skills collectively form a comprehensive skill library that enables the robot to handle various sub-tasks in complex navigation scenarios.
V Experiment
Figure 8: Box-pushing policy in different configurations.
(a) Different scenarios for real-world experiments.
(b) Box obstruction
(c) Box utilization: indoor
(d) Box utilization: outdoor
Figure 9: Visualization of real-world experiments. (a) shows various scenarios in real-world experiments. (b), (c) and (d) illustrate case studies for each scenario, in which the top-left subfigure shows the overall robot trajectory, while the subsequent key frames detail the specific process of reaching the navigation goals.
We also conduct real-world experiments to validate the performance of AINav in practical situations. We employ the Unitree Go2 as the robotic platform, which is equipped with an L1 Lidar for terrain information acquisition and an onboard RGB camera for visual detection. We evaluate the effectiveness of our method across three interactive navigation scenarios, as illustrated in Fig. 9(a). The first scenario involves an indoor obstacle obstruction setting, simulating a navigation task in a narrow corridor blocked by movable obstacles, where the robot needs to remove the box to reach the target. The second scenario, also set in a confined indoor corridor, featured a fixed $0.35$m-height platform obstructing the path. Since we assume the maximum climbing height for the robot is $0.3$ meters, the robot needs to reason how to traverse heights that exceed this capability limit by interacting with objects. The third scenario is box utilization in an outdoor environment, setting with fixed obstacles blocking the path and $0.45$m height, presenting a more difficult challenge with a different layout. Focusing primarily on the reasoning and planning capabilities of AINav, we utilize Apriltag detection to obtain accurate object descriptions, thereby mitigating the risk of task failure due to perception errors.
Fig. 9(b) demonstrates the experiment on obstacle obstruction. Initially, the robot detects a blocking box and navigates around it using local point cloud information and its navigation policy to approach the box. Subsequently, a pushing skill is employed to move the obstacle aside, creating a path to the target. Fig. 9(c) illustrates how the robot uses a box to overcome a higher obstacle. After visually detecting the box, the task planning module decides to push the box next to the high platform as a step, enabling the robot to climb over the challenging obstacle and reach the target. Fig. 9(d) shows the robot completing the interactive navigation task in the outdoor scenario, employing task decomposition and skill execution to achieve the target, overcoming challenges that traditional navigation methods could not address. Note that although the friction between indoor and outdoor scenarios is different, the robot can still push the box to the destination, demonstrating the robustness of the pre-trained pushing skill for interaction. These above real-world experiments demonstrate our method’s capability to complete navigation tasks in challenging environments. Further details are available in our supplementary video. We conduct $5$ repeated experiments in each scenarios, and the success rates are $100\%$ ($/5$ 5), $80\%$ ($4/5$), and $80\%$ ($4/5$), respectively, highlighting the robustness and superiority of AINav in various interactive tasks and both indoor and outdoor settings.
To evaluate the robustness of the pushing skill, we also carry out box-pushing tests with varying initial box positions in real-world scenarios, as shown in Fig. 8. We observe that the robot consistently pushes the box to the target position under different conditions, regardless of changes in the initial box placement. This stable performance across diverse configurations demonstrates the robustness of the pushing skills, facilitating effective interaction for navigation task completion.
VI Conclusion
We introduce an LLM-based adaptive interactive navigation approach using a quadruped robot, which proactively utilizes interactive objects to create feasible paths to reach originally unreachable goals. Our hierarchical framework leverages the LLM for task planning and employs pre-trained skills with RL for motion planning. The tree-structured task decomposition not only enhances the planning quality for long-horizon complex tasks but also allows for autonomous modifications in specific tree parts when new observations are obtained. Leveraging RL-driven pre-trained robust locomotion and interaction skills, the quadruped robot can manipulate the environment and create feasible paths to reach the navigation goal. Moreover, an adaptive replanning method is presented for intelligent understanding of incremental information, enabling robots to quickly adapt under partially observable situations. This instructive framework has the potential to enhance robotic navigation capabilities, offering an effective and computationally efficient solution for interactive navigation in challenging environments. Future directions encompass developing more generalizable interaction skills to accommodate broader tasks and scenarios, and achieving tighter integration between LLM/VLM and versatile control policies for more intelligent embodied planning and control.